OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x^3 + x^2 – 6x ?

OpenStudy (mertsj):

x^3+x^2-6x = 0 Factor out x x(x^2+x-6)=0 Factor the quadratic trinomial x((x+3)(x-2) = 0 Set all factors to 0 and solve x = 0, x+3=0 x-2=0 x=0,x=-3,x=2